### v2

# Multimodal Adversarial Defense for Vision-Language Models by Leveraging One-To-Many Relationships

Abstract
Pre-trained vision-language (VL) models are highly vulnerable to adversarial attacks. However, existing defense methods primarily focus on image classification, overlooking two key aspects of VL tasks: multimodal attacks, where both image and text can be perturbed, and the one-to-many relationship of images and texts, where a single image can correspond to multiple textual descriptions and vice versa (1:N and N:1). This work is the first to explore defense strategies against multimodal attacks in VL tasks, whereas prior VL defense methods focus on vision robustness. We propose multimodal adversarial training (MAT), which incorporates adversarial perturbations in both image and text modalities during training, significantly outperforming existing unimodal defenses. Furthermore, we discover that MAT is limited by deterministic one-to-one (1:1) image-text pairs in VL training data. To address this, we conduct a comprehensive study on leveraging one-to-many relationships to enhance robustness, investigating diverse augmentation techniques. Our analysis shows that, for a more effective defense, augmented image-text pairs should be well-aligned, diverse, yet avoid distribution shift—conditions overlooked by prior research. Our experiments show that MAT can effectively be applied to different VL models and tasks to improve adversarial robustness, outperforming previous efforts. Our code will be made public upon acceptance.
Vision-Language Model, Adversarial Robustness, Adversarial Defense
1 Introduction
Figure 1: Comparison of adversarial training (AT) methods for robust VL models. (a) Unimodal AT, such as TeCoA (Mao et al., 2022) and FARE (Schlarmann et al., 2024), robustifies a single modality via unimodal adversarial examples (AEs). However, it overlooks two key aspects of VL tasks: multimodal attacks, where attackers perturb both modalities, and one-to-many cross-modal alignment, where an image has multiple valid descriptions, and vice versa. (b) MAT addresses multimodal attacks by generating multimodal AEs during AT. (c) MAT+ further captures the inherent ambiguity in image-text relationships via one-to-many augmentations.
Vision-language (VL) tasks require modeling the relationships between image and text data. For example, image-text retrieval (ITR) aims to retrieve the most relevant text given an image query, and vice versa. Recent approaches leverage VL models pre-trained on large-scale image-text pairs, such as CLIP (Radford et al., 2021), which learns joint representations for images and texts via contrastive learning. However, recent studies revealed that these models are vulnerable to adversarial attacks (Zhang et al., 2022; Lu et al., 2023), which exploit nearly imperceptible perturbations in the input data. This poses significant risks of causing unintended consequences in real-world applications. For instance, in recommendation systems, attackers may manipulate images or their descriptions to alter the retrieval rankings, unfairly promoting or demoting specific entities. As VL models are deployed into society, understanding and mitigating their adversarial vulnerabilities has become increasingly urgent.
However, existing defense strategies for VL models (Mao et al., 2022; Wang et al., 2024b; Schlarmann et al., 2024) primarily focus on image attacks, e.g., for robust zero-shot image classification, leaving other VL tasks unexplored. This is a considerable oversight due to two critical aspects: (1) Multimodal manipulation: Attackers can manipulate both image and text modalities in VL tasks, while existing defense methods only consider image perturbations. Such an increased attack capability in VL tasks requires more complex defense strategies. (2) One-to-many (1:N) cross-modal alignment: Unlike classification, where text labels are simple and unambiguous (e.g., “a photo of a dog”), VL tasks involve sentence-level texts with a larger variation and ambiguity (e.g., the same image can be described as “a man with glasses is wearing a beer can crocheted hat” or “a man wears an orange hat and glasses”). This property makes achieving robust image-text alignment more complicated. Existing defense methods overlook these two key aspects, casting doubt on their effectiveness for VL tasks beyond zero-shot image classification.
To address this gap, we pioneer a study on defense strategies for VL models against multimodal attacks. Specifically, we study how to robustly fine-tune VL models for downstream VL tasks. First, we propose multimodal adversarial training (MAT), which incorporates both image-text perturbations during training. MAT significantly enhances robustness against multimodal attacks, demonstrating the necessity of defense methods tailored for multimodal attacks, an aspect overlooked by vision-only unimodal strategies (Mao et al., 2022; Schlarmann et al., 2024).
Furthermore, we argue that MAT’s performance is limited by the inadequate approximation of the real data distribution when using the deterministic (1:1) image-text pairs contained in the training data. To mitigate this, we investigate how to leverage the one-to-many (1:N) relationships inherent in VL data to enhance robustness. Inspired by works in cross-modal ambiguity modeling in ITR (Kim et al., 2023; Song & Soleymani, 2019), we explore various augmentation techniques to create diverse one-to-many (1:N) and many-to-one (N:1) image-text pairs. Our in-depth analysis reveals that augmentations are effective when image-text pairs are well-aligned and diverse, without causing distribution shift. For instance, text augmentations outperform image augmentations, whose higher dimensionality makes it difficult to avoid distribution shift. Additionally, cross-modal augmentations (e.g., $image\rightarrow text$) outperform intra-modal augmentations (e.g., $text\rightarrow text$) due to generating well-aligned image-text pairs. These are novel insights in the field of multimodal robustness, and complement the literature on unimodal adversarial training.
Our contributions are summarized as follows:
  • First defense strategy against multimodal attacks in VL models: We demonstrate that existing defense methods focused on vision are suboptimal for robust VL tasks and pioneer research in this new direction.
  • Proposed multimodal adversarial training (MAT): We largely improve robustness against multimodal attacks, highlighting the importance of considering multimodal perturbations in VL data.
  • Leveraging one-to-many relationships for robust VL tasks: We leverage one-to-many (1:N) image-text relationships via augmentations to enhance robustness, an aspect overlooked in unimodal adversarial training, which assumes a deterministic image-to-label mapping.
2 Related work
Adversarial attacks on vision-language models. Adversarial attacks on VL models are categorized into unimodal and multimodal. Unimodal attacks, such as gradient-based image attacks (Madry et al., 2017) and BERT-Attack for text (Li et al., 2020), perturb a single modality to mislead the models. In contrast, multimodal attacks, which perturb both image and text modalities, are significantly more effective (Zhang et al., 2022; Lu et al., 2023; Han et al., 2023; Wang et al., 2024a). However, developing defense strategies against multimodal attacks for VL tasks remains largely unexplored.
Adversarial defense for vision-language models. Existing defense strategies for VL models mainly focus on vision robustness, in which adversarial attacks perturb only the image. For example, Mao et al. (2022) and Wang et al. (2024b) approached zero-shot image classification on CLIP by proposing robust fine-tuning methods, which leverage unimodal adversarial training schemes to improve robustness. Schlarmann et al. (2024) also focused on image attacks only, and proposed a method for fine-tuning CLIP’s vision encoder to improve robustness in several VL tasks (i.e., image classification, image-text retrieval). Unlike the previous work, ours is the first to investigate adversarial defense strategies against multimodal attacks, where both image and text modalities are perturbed. Specifically, we propose a multimodal adversarial training strategy to enhance robustness against such attacks in VL models and tasks. Furthermore, we leverage the one-to-many (1:N) image-text relationships to further improve adversarial robustness.
Leveraging the one-to-many (1:N) nature of image-text. To tackle robustness in VL models, we take inspiration from current works that aim to model the ambiguity between image and text pairs; although a sentence can have multiple visual interpretations, normally only one is paired as the ground truth. Similarly, an image can be described using multiple different captions, but only one is considered as its pair. Since such 1:1 deterministic relationship is inconsistent with the 1:N nature of the data, existing works propose representing image-text samples as probabilistic embeddings (Chun et al., 2021; Chun, 2024), considering neighboring samples in the triplet loss (Thomas & Kovashka, 2020), and generating multiple and diverse representations for each image-text sample (Song & Soleymani, 2019; Kim et al., 2023). Among these solutions, the latter naturally fits the data augmentation strategy of adversarial training.
3 Preliminaries
Vision-language models. Many recent VL models (e.g., ALBEF (Li et al., 2021) and BLIP (Li et al., 2023)) are fundamentally built on CLIP, which learns joint image-text representations via large-scale image-text contrastive learning. CLIP consists of an image encoder $f_{\theta_{\text{I}}}:\mathbb{R}^{d_{\text{I}}}\to\mathbb{R}^{d_{\text{E}}}$ and a text encoder $f_{\theta_{\text{T}}}:\mathbb{R}^{d_{\text{T}}}\to\mathbb{R}^{d_{E}}$, where $\theta_{\text{I}}$ and $\theta_{\text{T}}$ are their parameters, $d_{\text{I}}$ and $d_{\text{T}}$ are input dimensions, and $d_{\text{E}}$ is the joint embedding dimension. Given an image $I\in\mathbb{R}^{d_{\text{I}}}$ and a text $T\in\mathbb{R}^{d_{\text{T}}}$, CLIP is trained to map them into a shared embedding space, maximizing the cosine similarity of image-text embeddings $S_{\theta_{\text{I},\text{T}}}(I,T)=\cos(f_{\theta_{\text{I}}}(I),f_{\theta_{\text{T}}}(T))$ for correct image-text pairs while minimizing it for incorrect pairs. CLIP optimizes the InfoNCE loss for a batch of $N$ image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{N}$ as:
$\mathcal{L}_{\text{CLIP-I}}(I,T)=-\sum_{i=1}^{N}\log\frac{\exp(S_{\theta_{\text{I},\text{T}}}(I_{i},T_{i})/\tau)}{\Sigma_{j=1}^{N}\exp(S_{\theta_{\text{I},\text{T}}}(I_{i},T_{j})/\tau)},$ (1)
where $\tau$ is the learnable temperature parameter. The overall loss is the average of the losses over images and texts, given by $\mathcal{L}_{\text{CLIP}}=(\mathcal{L}_{\text{CLIP-I}}+\mathcal{L}_{\text{CLIP-T}})/2$, where $\mathcal{L}_{\text{CLIP-T}}$ is the InfoNCE loss over texts.
Multimodal adversarial attacks. We aim to defend VL models against adversarial attacks, where both image and text modalities are perturbed. The objective of (untargeted) adversarial attacks on CLIP is to minimize the image-text similarity $S_{\theta_{I,T}}(I,T)$ for the correct image-text pairs $(I,T)$ to mislead the models’ predictions as:
$(I^{\prime},T^{\prime})$ $=(I+\delta_{\text{I}},T+\delta_{\text{T}})$ (2)
$=\operatorname*{arg\,min}_{\delta_{\text{I}}\in\Delta_{\text{I}},\delta_{\text{T}}\in\Delta_{\text{T}}}S_{\theta_{\text{I},\text{T}}}(I+\delta_{\text{I}},T+\delta_{\text{T}}),$
where $\delta_{\text{I}}$ and $\delta_{\text{T}}$ are image and text perturbations, and $\Delta_{\text{I}}$ and $\Delta_{\text{T}}$ define the set of allowed image and text perturbations. Image attacks maintain perceptual similarity between $I$ and $I^{\prime}$ typically using the $L_{p}$-norm. A common image attack strategy is projected gradient descent (PGD) (Madry et al., 2017), which iteratively updates $I^{\prime}$ by taking a small step in the direction of the gradient. Text attacks, such as BERT-Attack (Li et al., 2020), modify $N$ tokens in the text $T$ to maximize the divergence between $f_{\theta_{\text{T}}}(T)$ and $f_{\theta_{\text{T}}}(T^{\prime})$. Multimodal attacks perturb both image-text modalities to create $(I^{\prime},T^{\prime})$; Co-Attack (Zhang et al., 2022) perturbs each modality sequentially, first perturbing the image and then the text; SGA (Lu et al., 2023) enhances Co-Attack by considering the set-level interaction between multiple images and texts.
Adversarial training for image classification. Adversarial training (AT) (Madry et al., 2017) is a widely used defense against adversarial attacks, where a model is trained on adversarial examples. It solves a min-max optimization problem:
$\operatorname*{arg\,min}_{\theta_{\text{C}}}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left(\max_{\delta_{\text{I}}\in\Delta_{\text{I}}}\left[f_{\theta_{\text{C}}}(x+\delta_{\text{I}})\neq y\right]\right),$ (3)
where $\theta_{C}$ represents the parameters of image classifier $f_{\theta_{C}}$, and $(x,y)$ denotes an image and its corresponding class label drawn from the data distribution $\mathcal{D}$.
Based on this, to improve CLIP’s adversarial robustness in zero-shot image classification, TeCoA (Mao et al., 2022) adversarially fine-tunes CLIP’s image encoder by minimizing the contrastive loss between adversarial images and the text expression of the corresponding class (“a photo of {*}”):
$\operatorname*{arg\,min}_{\theta_{\text{I}}}\mathbb{E}_{(I,T)\sim\mathcal{D}}\left(\max_{\delta_{\text{I}}\in\Delta_{\text{I}}}\mathcal{L}_{\text{{CLIP-I}}}(I+\delta_{\text{I}},T)\right).$ (4)
However, TeCoA only defends against image adversarial attacks and does not consider the one-to-many (1:N) relationship in VL tasks.
4 Methodology
We propose Multimodal Adversarial Training (MAT) by addressing the limitations of existing unimodal AT (Fig. 1). First, we describe its practical optimization strategies via image-text contrastive learning applicable to VL models. Then, we introduce a data-driven approach that leverages the one-to-many relationship between images and texts to enhance robustness by more accurately approximating the true multimodal data distribution.
4.1 Problem setup
To defend against multimodal attacks, we introduce MAT, which perturbs both image and text modalities during AT. Following the theory of unimodal AT (Eq. 3), we formulate MAT objective as:
$\min_{\theta}\mathbb{E}_{(I,T)\sim\mathcal{D}^{\star}}\left(\max_{\begin{subarray}{c}\delta_{\text{I}}\in\Delta_{\text{I}},\\ \delta_{\text{T}}\in\Delta_{\text{T}}\end{subarray}}\mathcal{L}_{\text{VL}}(I+\delta_{\text{I}},T+\delta_{\text{T}})\right),$ (5)
where $\mathcal{D}^{\star}$ represents the true data distribution, $\theta$ represents the parameters of a VL model, and $\mathcal{L}_{\text{VL}}$ refers to the VL training loss (e.g., $\mathcal{L}_{\text{CLIP}}$ for CLIP).
4.2 Practical multimodal min-max optimization
4.2.1 Multimodal inner maximization
Directly optimizing the inner-maximization problem in Eq. 5 is highly complex, as it requires perturbing both modalities while accounting for their interactions. Nevertheless, in adversarial training, precise inner maximization is not always necessary; in fact, the basic PGD-AT defense method (Madry et al., 2017) uses approximation by training against first-order adversaries. Ensuring efficient optimization is equally important for feasible training time and scalability.
To this end, we propose two practical strategies for approximating the inner maximization: (1) Step-by-step perturbation: we generate adversarial examples $(I^{\prime},T^{\prime})$ by sequentially perturbing the text and image modalities, and (2) Loss function simplification: we replace the original loss function $\mathcal{L}_{\text{VL}}$ with a more efficient alternative when needed.
Adversarial text generation. First, we generate adversarial texts using a representative text attack, BERT-attack (Li et al., 2020). BERT-attack identifies critical words contributing to the loss and replaces them with semantically similar, grammatically correct alternatives to maximize the loss. However, this discrete optimization process is computationally expensive, and directly maximizing $\mathcal{L}_{\text{VL}}$ over multiple image-text pairs in a batch by repeatedly replacing words would be prohibitively time-consuming. To address this, we exploit the shared image-text alignment objective in VL models (e.g., CLIP, ALBEF, BLIP), and instead of direct loss maximization, we maximize the divergence between individual image-text embeddings as:
$T^{\prime}=\operatorname*{arg\,max}_{\delta_{\text{T}}\in\Delta_{\text{T}}}\|f_{\theta_{\text{T}}}(T+\delta_{\text{T}})-f_{\theta_{\text{I}}}(I)\|.$ (6)
Adversarial image generation. Next, given adversarial texts $T^{\prime}$, we generate adversarial images $I^{\prime}$ using PGD attack (Madry et al., 2017). Unlike the discrete optimization in text attacks, image attacks involve continuous optimization, where each update step requires a single back-propagation to maximize a loss function. Therefore, directly maximizing the loss $\mathcal{L}_{\text{VL}}$ is practically feasible. Nevertheless, for CLIP, instead of maximizing $\mathcal{L}_{\text{CLIP}}$, we adopt a more efficient approach by minimizing the cosine similarity between the image-text embeddings of $I^{\prime}$ and $T^{\prime}$ as:
$I^{\prime}=\operatorname*{arg\,max}_{\delta_{\text{I}}\in\Delta_{\text{I}}}-\frac{f_{\theta_{\text{I}}}(I+\delta_{\text{I}})\cdot f_{\theta_{\text{T}}}(T^{\prime})}{\|f_{\theta_{\text{I}}}(I+\delta_{\text{I}})\|\|f_{\theta_{\text{T}}}(T^{\prime})\|}.$ (7)
For ALBEF and BLIP, we directly maximize their downstream objective functions, as they involve advanced techniques that cannot be trivially simplified. Please see Appendix A.1 for details.
4.2.2 Multimodal outer minimization
To update the model parameters, we minimize the train loss function $\mathcal{L}_{\text{VL}}$ using the generated adversarial images and texts. For CLIP, we minimize the InfoNCE loss between the adversarial image $I^{\prime}$ and text $T^{\prime}$: $\min_{\theta}\mathcal{L}_{\text{CLIP}}(I^{\prime},T^{\prime})$. For ALBEF and BLIP, we fine-tune the models with their original downstream-task-specific objectives.
4.3 Leveraging one-to-many relationships for robustness generalization
While MAT addresses the multimodality of VL tasks, we hypothesize that robustness can be further improved by addressing also the one-to-many nature of image-text data. Thus, we propose Multimodal Augmented Adversarial Training (MAT+), which incorporates one-to-many augmentations to better approximate the true multimodal distribution of images and texts. Notably, this data-driven strategy for enhancing robustness is applicable to any VL model trained on image-text pairs.
4.3.1 Approximating the multimodal distribution by modeling ambiguity via augmentations
The true distribution of the data $\mathcal{D}^{\star}$ is inaccessible, but the dataset $\mathcal{\hat{D}}$ serves as an approximation, and its quality affects robustness significantly. In the field of image classification, Gowal et al. (2021) proved that adding image augmentations to the training data $\mathcal{\hat{D}}_{\text{tr}}$ produces a $\mathcal{\hat{D}}$ closer to $\mathcal{D}^{\star}$. This improves AT under the assumption of a deterministic closed-set image-to-label mapping. However, their theory does not consider the ambiguity of VL data, where a single image can have multiple valid descriptions and vice versa (Fig. 1). Thus, for VL data, multimodal AT with (1:1) image-text pairs only provides a limited $\mathcal{\hat{D}}$. To model this ambiguity, we generate diverse one-to-many (1:N) and many-to-one (N:1) image-text pairs through augmentations, improving the approximation of Eq. 5.
Modeling the ambiguity of text descriptions. Let us suppose a $\mathcal{\hat{D}}_{\text{tr}}$ consisting of deterministic (1:1) image-text pairs $(I,T)$. Relying solely on these pairs produces a weak $\mathcal{\hat{D}}$ because the original text description $T$ in $\mathcal{\hat{D}}_{\text{tr}}$ typically provides a partial or subjective depiction of the image $I$ (Fig. 1). In other words, $\mathcal{\hat{D}}_{\text{tr}}$ is typically constructed under the premise that $(I,T)$ are not ideal, but an approximation of the true image-to-text model $g_{\text{T}}^{\star}:\mathbb{R}^{d_{\text{I}}}\rightarrow\mathbb{R}^{d_{\text{T}}}$ (i.e., a “perfect” human annotator capable of generating exhaustive detailed descriptions). However, striving for a true $g_{\text{T}}^{\star}$ is neither desirable nor practical; language compresses visual information by focusing on human interests, naturally involving ambiguity, and approximating $g_{\text{T}}^{\star}$ deterministically would require a description detailed down to the individual pixel, which is linguistically unrealistic.
Thus, instead of relying on a single caption per image, we embrace text ambiguity by generating one-to-many (1:N) image-text pairs. In practice, given $\mathcal{\hat{D}}_{\text{tr}}$ with ground-truth pairs $(I,T)$, we can generate new aligned texts using $I$, $T$, or both. Consequently, Eq. 5 is approximated as:
$\min_{\theta}\mathbb{E}_{\begin{subarray}{c}(I,T)\sim\mathcal{\hat{D}}_{\text{tr}},\\ \phi\sim\Phi\end{subarray}}\left(\max_{\begin{subarray}{c}\delta_{\text{I}}\in\Delta_{\text{I}},\\ \delta_{\text{T}}\in\Delta_{\text{T}}\end{subarray}}\mathcal{L}_{\text{VL}}(I+\delta_{\text{I}},{\hat{g}}_{\text{T},\phi}(I,T)+\delta_{\text{T}})\right),$ (8)
where ${\hat{g}}_{\text{T},\phi}$ is the approximated image-to-text model, with $\phi$ capturing the inherent randomness (ambiguity) in generation, and $\Phi$ is the set of all possible random variables.
Modeling ambiguity of images. Similarly, a text description can correspond to multiple images, given the high dimensionality and degree-of-freedom of the image data space. For example, the description “a man with an orange hat” could be simultaneously paired with various images, such as a man with a green shirt or engaging in different activities (Fig. 1).
Thus, as in the text case, we introduce image augmentations to create many-to-one image-text pairs to approximate $\mathcal{D}^{\star}$. Formally, Eq. 5 is approximated as:
$\min_{\theta}\mathbb{E}_{\begin{subarray}{c}(I,T)\sim\mathcal{\hat{D}}_{\text{tr}},\\ \psi\sim\Psi\end{subarray}}\left(\max_{\begin{subarray}{c}\delta_{\text{I}}\in\Delta_{\text{I}},\\ \delta_{\text{T}}\in\Delta_{\text{T}}\end{subarray}}\mathcal{L}_{\text{VL}}({\hat{g}}_{\text{I},\psi}(I,T)+\delta_{\text{I}},T+\delta_{\text{T}})\right),$ (9)
where ${\hat{g}}_{\text{I},\psi}$ is the approximated text-to-image model, with $\psi$ capturing the inherent randomness, and $\Psi$ denoting the set of all possible random variables.
4.3.2 Conditions for effective augmentations
Not all $\Phi$ and $\Psi$ produce a model that can generate valid image-text pairs ×$\mathcal{I}\times\mathcal{T}\subseteq\mathbb{R}^{d_{\text{I}}}\times\mathbb{R}^{d_{\text{T}}}$, that is, paired data that effectively refines $\mathcal{\hat{D}}$ towards $\mathcal{D}^{\star}$. We can develop the theory of Gowal et al. (2021) for image classification to address multimodal VL augmentations. Let $p$ be the probability measure corresponding to $\mathcal{D}^{\star}$, for which every valid pair $(I,T)\in\mathcal{I}\times\mathcal{T}$ has non-zero probability: $p(I,T)>0$. Similarly, let $\hat{p}$ be the probability measure corresponding to $\mathcal{\hat{D}}$. The sufficient condition for $\mathcal{\hat{D}}$ to be an effective approximation is as follows:
Condition 1.
(Accurate approximated distribution) The approximated data distribution $\mathcal{\hat{D}}$ and true data distribution $\mathcal{D}^{\star}$ must be equivalent: $\forall(I,T)\in\mathcal{I}\times\mathcal{T},p(I,T)=\hat{p}(I,T).$
Ineffective one-to-many augmentations. Let us denote augmentations for image-text pairs $(I,T)\in\mathcal{\hat{D}}_{\text{tr}}$ as $I_{\text{aug}}=\hat{g}_{\text{I},\psi}(I,T)$ and $T_{\text{aug}}=\hat{g}_{\text{T},\phi}(I,T)$. Then, when adding an augmented pair $(I_{\text{aug}},T)$ or $(I,T_{\text{aug}})$ to $\mathcal{\hat{D}}_{\text{tr}}$, we define three cases that result into ineffective augmentations, violating Cond. 1.
Case 1.
(Semantic mismatch) If augmented image-text pairs are not semantically aligned, $\hat{p}$ diverges from $p$ by generating data in regions of zero probability; that is, $\hat{p}(I_{\text{aug}},T)\neq p(I_{\text{aug}},T)=0$ or $\hat{p}(I,T_{\text{aug}})\neq p(I,T_{\text{aug}})=0$, meaning that the distributions $\mathcal{D}^{\star}$ and $\mathcal{\hat{D}}$ differ.
Case 2.
(Limited diversity) Even if augmented image-text pairs are semantically aligned, trivial augmentations $I\simeq I_{\text{aug}}$ or $T\simeq T_{\text{aug}}$ do not provide enough variation to refine $\mathcal{\hat{D}}$ and capture the inherent ambiguity.
Case 3.
(Distribution shift) Even if augmented image-text pairs are aligned, excessive augmentations cause distribution shifts. This also applies to the unimodal case (Gowal et al., 2021).
To make sure our augmentations are effective, we introduce three measurable properties as follows:
  • Alignment: The semantic similarity between augmented image-text pairs: $S_{\theta_{\text{I},\text{T}}}(I_{\text{aug}},T)$ or $S_{\theta_{\text{I},\text{T}}}(I,T_{\text{aug}})$.
  • Diversity: The Kullback-Leibler (KL) divergence between the original and augmented samples: $d_{\text{KL}}(I,I_{\text{aug}})$ or $d_{\text{KL}}(T,T_{\text{aug}})$.
  • Distribution gap: The Fréchet distance (Fréchet, 1936) between the distributions of the original and augmented samples, $d_{\text{F}}(\mathcal{\hat{D}}_{\text{tr}},\mathcal{\hat{D}}_{\text{aug}})$, where $\mathcal{\hat{D}}_{\text{aug}}=\{(I_{\text{aug}},T),(I,T_{\text{aug}})\}^{N}_{i=1}$. This metric, widely used for evaluating generative models (e.g., FID (Heusel et al., 2017)), quantifies the distribution gap.
High alignment reduces semantic mismatches (Case 1), sufficient diversity captures image-text ambiguity, preventing Case 2, and minimal distribution gap avoids generating out-of-distribution samples, preventing Case 3.
Our results in Section 6 show that cross-modal augmentations (image $\leftrightarrow$ text) achieve better alignment than intra-modal augmentations (image $\to$ image, text $\to$ text), which cannot capture image-text interactions. Moreover, achieving both sufficient diversity and minimal distribution gap is harder to achieve in image augmentations due to the high dimensionality and degree-of-freedom of the image space.
5 Experimental settings
Our evaluation is focused mainly on improving robustness for image text retrieval (ITR), the most fundamental VL task. For this, we apply MAT to three representative VL models: CLIP, ALBEF, and BLIP. Finally, we also evaluate additional VL tasks, i.e., ALBEF on visual grounding (VG) and BLIP on image captioning (IC).
Method Img aug. Text. aug. Clean Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 TR@5 IR@1 IR@5
Finetune 92.1 77.2 11.0 6.7 0.6 2.0 0.6 2.6
[…]
TeCoA-ITR 83.1 68.2 52.9 31.9 27.5 52.4 17.6 40.1
(ours) MAT 83.7 67.5 56.9 37.1 37.5 65.1 24.8 49.7
[…]
I2T(div-Caps) 84.7 $\uparrow$1.0 69.2 $\uparrow$1.7 60.3 $\uparrow$3.4 39.6 $\uparrow$2.6 40.3 $\uparrow$2.8 66.4 $\uparrow$1.3 27.8 $\uparrow$3.0 54.3 $\uparrow$4.5
I2T(Human) 85.7 $\uparrow$2.0 71.9 $\uparrow$4.4 63.3 $\uparrow$6.4 44.0 $\uparrow$6.9 45.6 $\uparrow$8.1 70.4 $\uparrow$5.3 32.2 $\uparrow$7.4 58.4 $\uparrow$8.6
Basic(RandAug) 84.1 $\uparrow$0.4 67.1 $\downarrow$0.4 57.8 $\uparrow$0.9 37.3 $\uparrow$0.2 35.6 $\downarrow$1.9 64.3 $\downarrow$0.8 24.4 $\downarrow$0.4 50.2 $\uparrow$0.4
T2I(SD) 83.3 $\downarrow$0.4 68.4 $\uparrow$0.9 57.2 $\uparrow$0.3 37.8 $\uparrow$0.7 37.9 $\uparrow$0.4 63.4 $\downarrow$1.7 25.2 $\uparrow$0.4 50.1 $\uparrow$0.3
T-I2I(SD) 83.8 $\uparrow$0.1 69.1 $\uparrow$1.6 59.7 $\uparrow$2.8 38.0 $\uparrow$1.0 39.3 $\uparrow$1.8 66.1 $\uparrow$1.0 25.8 $\uparrow$1.0 51.2 $\uparrow$1.5
Table 1: Comparison of CLIP trained on the Flickr30k dataset for image-text retrieval (ITR) under the no-attack scenario (Clean) and multimodal attacks (Co-Attack, SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Method Img aug. Text. aug. Clean Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 TR@5 IR@1 IR@5
Finetune 66.6 50.1 2.9 1.8 0.1 0.5 0.1 0.6
[…]
TeCoA-ITR 58.0 41.6 21.3 12.5 9.6 24.4 6.2 17.6
(ours) MAT 55.8 40.7 31.4 19.5 17.7 39.1 12.3 30.8
[…]
I2T(div-Caps) 56.7$\uparrow$0.9 39.9$\downarrow$0.8 30.3$\downarrow$1.1 19.4$\downarrow$0.2 18.9$\uparrow$1.2 40.2$\uparrow$1.1 12.5$\uparrow$0.2 30.2$\downarrow$0.6
I2T(Human) 58.9$\uparrow$3.1 43.1$\uparrow$2.4 33.6$\uparrow$2.2 21.9$\uparrow$2.3 21.3$\uparrow$3.6 43.2$\uparrow$4.1 14.5$\uparrow$2.2 34.6$\uparrow$3.8
Basic(RandAug) 55.9$\uparrow$0.1 40.7$\uparrow$0.0 30.7$\downarrow$0.7 19.6$\uparrow$0.1 18.3$\uparrow$0.6 40.1$\uparrow$1.0 12.6$\uparrow$0.3 31.3$\uparrow$0.6
T2I(SD) 54.2$\downarrow$1.6 38.3$\downarrow$2.4 28.5$\downarrow$2.9 18.1$\downarrow$1.5 17.2$\downarrow$0.5 37.5$\downarrow$1.6 11.7$\downarrow$0.6 29.9$\downarrow$0.9
T-I2I(SD) 57.1$\uparrow$1.3 41.7$\uparrow$1.0 31.3$\downarrow$0.0 20.2 $\uparrow$0.7 18.8 $\uparrow$1.1 39.7 $\uparrow$0.6 12.6 $\uparrow$0.2 31.0 $\uparrow$0.3
Table 2: Comparison of CLIP trained on the COCO dataset for image-text retrieval (ITR) under the no-attack scenario (Clean) and multimodal attacks (Co-Attack, SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Datasets. We use the commonly used Flickr30k (Plummer et al., 2015) and the COCO (Chen et al., 2015) datasets for ITR. While these datasets contain five captions per image, our baseline training uses 1:1 image-text pairs (taking the first annotated caption), reflecting the typical setup when fine-tuning with real-world data, such as data collected from the internet, where 1:1 pairings are more common. For VG, we use RefCOCO+ (Kazemzadeh et al., 2014), and for IC, we use COCO. Please see Appendix A for details.
Evaluation metrics. We use the recall@k (R@k) for evaluating ITR, including both image-to-text and text-to-image retrieval. For VG we use the accuracy in localizing the regions corresponding to the text descriptions, while IC evaluates the quality of generated captions using diverse metrics. Details are in Appendix A.
Adversarial attacks. We evaluate our defense methods against multimodal adversarial attacks, Co-Attack (Zhang et al., 2022) and SGA (Lu et al., 2023). The perturbation constraints are $\epsilon_{I}=2/255$ ($L_{\infty}$-norm) for image and one token for text, following (Zhang et al., 2022; Lu et al., 2023). Additional results on unimodal attacks, including PGD and BERT-Attack, are in Appendix C.
Baseline methods. Since there do not exist defense methods tailored for multimodal attacks, we compare our method with the unimodal AT methods closest to our setting proposed for CLIP:
  • FARE (Schlarmann et al., 2024): An unsupervised adversarial fine-tuning scheme for CLIP, which focuses on obtaining a robust CLIP vision encoder. Since it is unsupervised, we can apply it to our ITR setting as-is.
  • TeCoA-ITR: Since the original TeCoA (Mao et al., 2022) is supervised, we extend it to be used in ITR. While TeCoA fine-tunes the vision encoder for image classification, TeCoA-ITR fine-tunes both vision and text encoders using Eq. 7 to generate adversarial images.
Training details. We fine-tune the pre-trained CLIP-ViT-B/16, ALBEF-14M, and BLIP w/ ViT-B models using MAT. For this, we generate adversarial images using 2-step-PGD (perturbation size of 2/255 in $l_{\infty}$-norm), and adversarial texts using BERT-attack (perturbation size of 1 token). Please see the detailed hyperparameters in Appendix A.4.
Method Img aug. Text aug. Flickr30k COCO
Clean SGA Clean SGA
TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1
Finetune 89.5 77.7 2.5 1.3 69.9 53.6 1.0 0.7
[…]
(ours) MAT 82.0 66.3 47.1 32.9 63.9 46.2 31.2 21.2
[…]
I2T(div-Caps) 85.6 $\uparrow$3.6 71.0 $\uparrow$4.8 48.8 $\uparrow$1.7 35.0 $\uparrow$2.1 66.0 $\uparrow$2.0 49.9 $\uparrow$3.7 35.5 $\uparrow$4.3 20.3 $\downarrow$0.9
I2T(Human) 85.8 $\uparrow$3.8 72.8 $\uparrow$6.5 52.9 $\uparrow$5.8 38.8 $\uparrow$5.9 68.5 $\uparrow$4.5 49.1 $\uparrow$3.0 36.2 $\uparrow$4.9 23.5 $\uparrow$2.2
Basic(RandAug) 82.2 $\uparrow$0.2 67.2 $\uparrow$0.9 48.3 $\uparrow$1.2 33.4 $\uparrow$0.5 63.1 $\downarrow$0.9 48.3 $\uparrow$2.1 30.3 $\downarrow$1.0 21.7 $\uparrow$0.4
T2I(SD) 83.7 $\uparrow$1.7 68.3 $\uparrow$2.1 52.0 $\uparrow$4.9 36.2 $\uparrow$3.3 61.5 $\downarrow$2.4 46.0 $\downarrow$0.1 25.3 $\downarrow$5.9 18.6 $\downarrow$2.6
T-I2I(SD) 85.1 $\uparrow$3.1 69.8 $\uparrow$3.6 55.2 $\uparrow$8.1 37.6 $\uparrow$4.8 64.9 $\uparrow$0.9 47.1 $\uparrow$1.0 33.2 $\uparrow$2.0 22.4 $\uparrow$1.2
Table 3: Comparison of ALBEF trained on Flickr30k and COCO datasets for image-text retrieval (ITR) under the no-attack scenario (Clean) and a multimodal attack (SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Method Aug. Clean SGA
TR@1 IR@1 TR@1 IR@1
Finetune 72.9 57.5 1.2 1.1
[…]
(ours) MAT 66.9 49.9 31.3 21.0
[…]
T-I2I(SD) 68.2 $\uparrow$1.3 50.5 $\uparrow$0.6 33.5 $\uparrow$2.2 22.9 $\uparrow$1.9
Table 4: Comparison of BLIP trained on COCO for ITR under the no-attack scenario (Clean) and the SGA attack, reporting R@k.
Augmentation strategies. We consider two types of augmentations: intra-modal and cross-modal. Intra-modal augmentation enhances data points without considering image-text interactions (text $\rightarrow$ text, image $\rightarrow$ image), while cross-modal augmentation enhances data points by leveraging the other modality (image $\leftrightarrow$ text). Specifically, we explore the following augmentation techniques:
  • Text augmentations:
    • Intra-modal: EDA (Wei & Zou, 2019) for basic word-level edits. LLM-based rewriting (Fan et al., 2024) is shown in Appendix C.
    • Cross-modal: Image-to-text (I2T) generation using diverse prompts with InternVL (Chen et al., 2024) (“I2T(div-Caps)”), and human-generated captions from the remaining 4 annotations of Flickr30k and COCO (“I2T(Human)”).
  • Image augmentations:
    • Intra-modal: RandAug (Cubuk et al., 2020), randomly applying affine transformations and color distortions.
    • Cross-modal: Stable Diffusion (SD) for text-to-image (T2I) (Rombach et al., 2022) and text-guided image-to-image (T-I2I) (Meng et al., 2021).
Each original data point is augmented four times, leading to a $\times$5 expansion, except for T-I2I(SD), where using only two augmentations yielded better results. Please see Appendix A.3 for the detailed augmentation settings.
6 Results
6.1 Image-text retrieval (ITR)
Table 1 and 2 present the results for ITR using CLIP trained on Flickr30k and COCO, respectively. Table 3 presents the results for ALBEF trained on the same datasets, while Tab. 4 evaluates BLIP on COCO.
6.1.1 Effectiveness of Multimodal Adversarial Training
By introducing multimodal adversarial perturbations, MAT consistently achieves significantly greater robustness against multimodal attacks than the unimodal AT methods, FARE and TeCoA-ITR, which focus solely on image perturbations during AT. The improvements are substantial and consistent for CLIP on Flickr30k (Tab. 1) and COCO (Tab. 2), as well as ALBEF on both datasets (Tab. 3). These results highlight the necessity of defense strategies tailored for multimodal perturbations.
Figure 2: The relationships between the three properties of augmentations, (i) alignment, (ii) diversity, and (iii) distribution gap, versus Robust Accuracy against multimodal attack (the overall average of IR@1/TR@1 for CLIP/ALBEF on Flickr/COCO).
6.1.2 Effectiveness of one-to-many augmentations
We observe that effective augmentations, such as I2T(Human), consistently enhance robustness, whereas some augmentations, such as T2I(SD), do not.
To identify key factors for effective augmentation, we analyze alignment, diversity, and distribution gap based on our hypothesis for approximating the true image-text distribution (Sec.4.3.2). Figure 2 presents these scores, computed in CLIP’s embedding space. Alignment measures the cosine similarity between augmented image-text pairs, diversity quantifies the average KL-divergence between each original and augmented sample, and distribution gap measures the overall distance between the distributions of original and augmented samples. Figure 2 reveals that all properties, high alignment, high diversity, and small distribution gap, is crucial for effective augmentations.
Cross-modal augmentation outperforms intra-modal augmentation due to higher image-text alignment. We observe that cross-modal augmentations can result in better robustness than intra-modal augmentations. This can be attributed to cross-modal augmentations taking into account the other modality, generating well-aligned image-text pairs. For example, the alignment scores of RandAug and EDA are lower than those of the original pairs (Fig. 2), indicating semantic mismatch (Case. 1). Intra-modal augmentations struggle to balance diversity and alignment: increasing diversity without considering image-text relationships disrupts alignment. This is evident in Fig. 2, where intra-modal augmentations appear in the lower-left, achieving only one of alignment or diversity.
Achieving sufficient diversity while keeping the distribution gap minimal is crucial. Figure 2 indicates that even if augmentations are well-aligned as well as sufficiently diverse to expand the training data, they do not enhance robustness when there is a large distribution gap from the original samples. For example, T2I(SD) introduces a good diversity but also introduces a large distribution gap, resulting in minimal robustness improvement.
Appendix B provides qualitative examples of the augmentations generated for each technique.
(a) T-I2I(SD)
(b) I2T(Human)
Figure 3: Analysis of the number of augmentations and robustness against SGA attacks for: (a) image augmentations using T-I2I(SD), and (b) text augmentations using I2T(Human).
Image augmentations are challenging. While both image and text augmentations proved effective, the latter performed comparatively better. The reason lies in the high dimensionality and degree of freedom of the image space, which makes it harder to generate augmentations that are both diverse and stay within the distribution. In contrast, the lower-dimensional, structured nature of text makes it more amenable to controlled augmentation. Figure 2 shows that text augmentations can introduce diversity while maintaining a small distribution gap, whereas image augmentations cannot. Additionally, Fig. 3 analyzes the impact of the number of augmentations on multimodal robustness. We observed that increasing text augmentations with I2T(Human) improves robustness, as more captions better approximate the data distribution, however, image augmentations using T-I2I(SD) enhance robustness up to two additional augmentations but can degrade it beyond that due to distribution shift. Therefore, we believe that developing image augmentations with sufficient diversity while minimizing the distribution gap is a promising direction for future research.
Many-to-many augmentations (N:N). Appendix C shows that combining image and text augmentations (N:N) did not enhance performance over text-only augmentations. With two augmented images and four texts, original pairs constitute only a 12.5% of the data, which makes (N:N) augmentations prone to distorting the data distribution if not designed carefully. Thus, while theoretically promising, many-to-many augmentations require a dedicated methodology, which falls out of the scope of this work.
Method Aug. SGA
Val Test-A Test-B
Fine-tune 32.9 36.3 29.6
[…]
(ours) MAT 40.4 44.2 35.1
[…]
I2T(Human) 42.0 $\uparrow$1.6 46.5 $\uparrow$2.3 35.7 $\uparrow$0.6
T-I2I(SD)))) 41.1 $\uparrow$0.7 46.0 $\uparrow$1.8 34.8 $\downarrow$0.3
Table 5: Accuracy comparison using ALBEF on the RefCOCO+ dataset for visual grounding.
Method Aug. SGA
BLEU-4 ROUGE CIDEr
Finetune 11.1 35.6 35.5
[…]
(ours) MAT 21.8 46.8 74.4
[…]
Table 6: Generated text quality comparison using BLIP evaluated on the COCO dataset for IC.
6.2 Additional tasks
Visual grounding (VG). We evaluate MAT in VG using the model and dataset originally proposed for this task, ALBEF and RefCOCO+ (Tab. 5). MAT proves to be the most effective in this task as well, while MAT+ further improves robustness. This demonstrates the effectiveness of one-to-many augmentations in VL tasks beyond ITR.
Image captioning (IC). We verify the effectiveness of one-to-many augmentations in IC using the model and dataset originally proposed for this task, BLIP and COCO (Tab. 6). Since IC involves only image inputs, MAT does not surpass FARE, the unimodal AT method tailored for image attacks. Nevertheless, we demonstrate that augmentations enhance robustness in this task, with MAT+ outperforming FARE.
7 Conclusions
This is the first work to study adversarial defense strategies for vision-language (VL) models against multimodal adversarial attacks in VL tasks. Existing defense methods are ineffective against multimodal attacks, as they focus on defending against image-only attacks, and also ignore the one-to-many nature of images and texts. To address this, we proposed MAT, a novel defense framework that leverages multimodal perturbations and one-to-many augmentations. Our comprehensive analysis reveals that our framework significantly improves adversarial robustness against multimodal attacks and that augmentations are most effective when well-aligned, diverse, and avoid distribution shift; a setting that previous literature on unimodal adversarial defense did not explore. Our novel perspective identified several challenges, such as the difficulty of generating image augmentations without distorting the original joint distribution of images and texts, opening a promising path for future research.
Appendix A Experimental Details
A.1 MAT’s Inner Maximization Strategy
Model VL Task Inner maximization Outer minimization
Text Adv. Generation Image Adv. Generation
CLIP Image-Text Retrieval $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\min_{\delta_{I}}\cos(f_{\theta_{\text{I}}}(I+\delta_{I}),f_{\theta_{\text{T}}}(T))$ $\mathcal{L}_{\text{CLIP}}$
ALBEF Image-Text Retrieval $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
Visual Grounding $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
BLIP Image-Text Retrieval $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
Image Captioning $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{LM}}$ $\mathcal{L}_{\text{LM}}$
Table 7: Summary of MAT’s strategies for each model and downstream task.
In the main text, we proposed MAT, a multimodal defense framework and practical optimization strategies. In our experiments, we adaptively modify MAT’s inner maximization strategy for different downstream tasks, as summarized in Table 7. We provide detailed explanations below.
A.1.1 Image-Text Retrieval (ITR)
MAT for CLIP generates text adversarial examples by maximizing the divergence between image-text embeddings, using BERT-attack (Li et al., 2020), while image adversarial examples are generated by minimizing their cosine similarity, following image attacks in Co-Attack (Zhang et al., 2022) and SGA attack (Lu et al., 2023).
The ALBEF and BLIP model architectures use a similar approach for ITR but include additional components, such as an ITM module (Image-Text Matching) and multimodal encoders. ITC loss ($\mathcal{L}_{\text{ITC}}$) is the image-text contrastive loss, similar to $\mathcal{L}_{\text{CLIP}}$, with the primary difference being whether momentum encoders are used to store previously seen representations. ITM loss ($\mathcal{L}_{\text{ITM}}$) predicts whether an image-text pair is positive (matched) or negative (not matched), which is predicted by a multimodal encoder in ALBEF and BLIP in addition to their unimodal image and text encoders. Please refer to original papers (Li et al., 2021, 2023) for more details. Given the added complexity in ALBEF and BLIP, MAT for these models is designed as follows:
  • Text attack: Same strategy as training CLIP with MAT.
  • Image attack: Directly maximize the downstream-task specific fine-tuning objective.
A.1.2 Visual Grounding (VG)
In our experiment, we evaluated the effectiveness of MAT on visual grounding (VG) using ALBEF. Since the fine-tuning objective for VG is the same as for ITR, the inner maximization strategy of MAT is identical to MAT for ITR.
A.1.3 Image Captioning
In our experiment, we evaluated the effectiveness of MAT on image captioning using BLIP. The fine-tuning objective for image captioning includes ITC loss and language modeling (LM) loss, and we maximize the sum of these objectives for perturbing images.
A.2 Evaluation Settings
A.2.1 Image-Text Retrieval (ITR)
Dataset. We use Flickr30k, which consists of image-text pairs with a train/test split of 29,000/1,000 images, and COCO, which consists of image-text pairs with a train/val/test split of 113,287/5,000/5,000 images. We use the default split for training and testing. While each image has five captions, our baseline training approach uses 1:1 image-text pairs, reflecting the practical setting for fine-tuning VL models.
Evaluation metric. We use R@k, which measures the recall of the correct image or text among the top-k retrieved candidates. These metrics assess the model’s ability to correctly retrieve relevant images or texts when given a query.
A.2.2 Visual Grounding (VG)
Dataset. For training, we use MSCOCO’s training set, following the ITR setting. For testing VG tasks, we use RefCOCO+, which includes text descriptions of objects in images along with their corresponding bounding boxes. RefCOCO+ contains 141,564 expressions for 19,992 images from the COCO training set. This dataset was collected interactively through a two-player game (Kazemzadeh et al., 2014). It consists of three test sets, Val, TestA, and TestB. Test Set A contains objects sampled randomly from the entire dataset. Test Set B contains objects sampled from the most frequently occurring object categories in the dataset. Test Set C contains objects sampled from images that contain at least 2 objects of the same category.
Evaluation metric. VG aims to localize the region in an image that corresponds to a specific text description. Following Li et al. (2021), we extend Grad-CAM (Selvaraju et al., 2017) to acquire heatmaps, and use them to rank the detected proposals provided by Yu et al. (2018). We measure the accuracy of the attention maps with the IoU threshold being 0.5.
A.2.3 Image Captioning
Dataset. We use the MSCOCO dataset for both training and testing. For training data, we follow the ITR setting.
Evaluation metric. We use various evaluation metrics to measure the quality of the generated captions, including BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE (Lin & Och, 2004), CIDEr (Vedantam et al., 2015), and SPICE (Anderson et al., 2016).
A.3 Augmentation techniques
A.3.1 Text augmentations
EDA (Easy Data Augmentation). EDA randomly selects words in the text and performs the following operations: synonym replacement, random insertion, random swap, or random deletion. We use the official implementation  [^1] https://github.com/jasonwei20/eda_nlp. The hyperparameter $\alpha$ controls the strength of the augmentation, where $\alpha$ determines the probability of each word being augmented. We use $\alpha=0.3$ for all experiments.
LangRW (Language rewrite). Language rewrite (LangRW) (Fan et al., 2024) is a method that rewrites the text data to improve the robustness of the model, using a generative natural language processing model, such as Llama (Touvron et al., 2023). We used Llama-2-7B  [^2] https://huggingface.co/meta-llama/Llama-2-7b. In our work, we used slightly modified prompts from the original work to simultaneously generate four captions per image. Given an original caption $T$, the prompt for generating additional captions are as follows:
     Rewrite image captions in 4 different  ways.    {coco caption 1 for image i}  => {coco caption 2 for image i}  => {coco caption 3 for image i}  => {coco caption 4 for image i}  => {coco caption 5 for image i}    {coco caption 1 for image j}  => {coco caption 2 for image j}  => {coco caption 3 for image j}  => {coco caption 4 for image j}  => {coco caption 5 for image j}    {coco caption 1 for image k}  => {coco caption 2 for image k}  => {coco caption 3 for image k}  => {coco caption 4 for image k}  => {coco caption 5 for image k}    {original caption to be rewritten}  =>     where the coco captions are randomly sampled from the original captions from the COCO dataset (Chen et al., 2015).
I2T(div-Caps) using div-Caps. div-Caps (Chen et al., 2024) is a latest vision-language multimodal model. We use the div-Caps-2.5-2B model. We generate four captions per image using the following four prompts to ensure diversity:
  • Details: “Describe the image in detail.”
  • MainObj: “Describe only the one main object in the image, do not say anything about the other objects or background.”
  • Background: “Describe only the background of this image, do not say anything about the foreground objects.”
  • Style: “Describe the style or your feelings about this image, do not say anything about the objects in the image.”
The lower alignment of div-Caps-generated captions compared to original image-text pairs (Fig. 2) is due to the fact that captions generated with Background and Style are less aligned with images than the full captions.
Human. Human augmentation is a method that generates additional captions by human annotators. Since we use 1:1 image-text pairs for training as default, we used the rest of the original captions included in Flickr30k and COCO datasets as additional captions for each image.
A.3.2 Image augmentations
RandAug (Random Augmentation). RandAug (Cubuk et al., 2020) is an image augmentation method that applies a series of random transformations to the image. We used the codes from ALBEF (Li et al., 2021)  [^3] https://github.com/salesforce/ALBEF. We set the number of operations to 2 and the magnitude to 5 for all experiments.
Stable Diffusion (SD); Text-to-Image. We used SD-v2.1  [^4] https://huggingface.co/stabilityai/stable-diffusion-2-1-base for text-to-image augmentations, using Huggingface’s default hyperparameters.
[…]
A.4 Hyperparameter Details
We fine-tune the pre-trained CLIP-ViT-B/16, ALBEF-14M, and BLIP w/ ViT-B. We generate adversarial images using 2-step-PGD with a step size of $1.0/255$ with the perturbation bound $\epsilon_{\text{I}}=2.0/255$. For adversarial texts, we generate with BERT-attack, which replaces a word from top-k candidates with $k=10$. We train for 5,000 steps on Flickr30k and 10,000 steps on COCO, with a batch size of 128. For CLIP, we train using the SGD optimizer, while for ALBEF and BLIP, we use the AdamW optimizer. All models are trained with cosine learning rate scheduling, a learning rate of 0.0001, and a weight decay of 0.0001.
Appendix B Visualization of Augmentations
This section visualizes each augmentation technique. Figure 4 and Fig. 5 show image augmentations, and Fig. 6 and Fig. 7 visualize text augmentations. These visualizations help in understanding image-text alignment, augmentation diversity, and the distribution of augmentations. For example, Fig.4 and Fig.5 show that while I2T(SD) generates diverse images, they differ significantly from the original images and often appear somewhat synthetic, potentially causing a distribution shift. Additionally, Fig. 6 and Fig. 7 show that Basic(EDA), which is a basic word-level augmentation, can disrupt the image-text alignment.
Figure 4: Visualization of image augmentations (1).
Figure 5: Visualization of image augmentations (2).
Figure 6: Visualization of text augmentations (1).
Figure 7: Visualization of text augmentations (2).
Appendix C Additional Results
C.1 Image-text retrieval (ITR)
In this section, we present comprehensive results for image-text retrieval (ITR), evaluating defense methods against unimodal attacks (PGD (Madry et al., 2017) for image attack and BERT-attack (Li et al., 2020)) as well as against multimodal attacks (Co-Attack (Zhang et al., 2022) and SGA (Lu et al., 2023)). Table 8 and Table 9 present the results of CLIP trained on Flickr30k and COCO, respectively. Table 10 and Table 11 show the results of ALBEF trained on Flickr30k and COCO, while Table 12 presents the results of BLIP trained on COCO.
First, the results demonstrate that MAT consistently outperforms baseline unimodal AT methods in terms of multimodal robustness, which is crucial given that multimodal attacks are significantly stronger than unimodal attacks in VL tasks. The robustness against image attacks (PGD) is generally similar between image-only defenses (FARE or TeCoA-ITR) and MAT, which aligns with expectations.
Second, the results indicate that effective augmentations, such as I2T(Human) for text augmentation and T-I2I(SD) for image augmentation, consistently improve multimodal robustness. The important factors that determines the effectiveness of augmentations are described in Sec. 6: both I2T(Human) and T-I2I(SD) augment well-aligned image-text pairs with sufficient diversity, while keeping their distribution close to the original samples.
In addition to the main paper, we present results on many-to-many augmentation in Tab. 9, where we combine image and text augmentations—specifically, T-I2I(SD) for image augmentation and I2T(Human) for text augmentation—–aiming at enhancing robustness. However, our results show that this combination does not provide further improvements over using I2T(Human) alone, suggesting that the added image augmentation does not contribute to additional robustness. A possible reason for this is that T-I2I(SD) introduces a slight distribution shift, which may offset its benefits. These findings indicate that while many-to-many augmentations have theoretical potential, they require augmentation techniques that introduce sufficient diversity while maintaining consistency with the original data distribution.
Method Img aug. Text. aug. Unimodal attacks Multimodal attacks
Clean PGD (image attack) BERT (text attack) Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1
Finetune 92.1 77.2 11.9 10.1 75.4 53.1 11.0 6.7 0.6 0.6
FARE 75.9 61.0 69.7 55.1 53.2 40.2 41.8 30.7 27.1 21.0
TeCoA-ITR 83.1 68.2 77.7 61.9 64.7 42.7 52.9 31.9 27.5 17.6
(ours) MAT 83.7 67.5 77.4 61.4 72.2 51.1 56.9 37.1 37.5 24.8
[…]
T2T(LangRW) 80.9 $\downarrow$2.8 67.0 $\downarrow$0.5 74.2 $\downarrow$3.2 60.1 $\downarrow$1.3 72.6 $\uparrow$0.4 51.3 $\uparrow$0.3 58.9 $\uparrow$2.0 38.9 $\uparrow$1.9 40.0 $\uparrow$2.5 27.4 $\uparrow$2.6
I2T(div-Caps) 84.7 $\uparrow$1.0 69.2 $\uparrow$1.7 79.0 $\uparrow$1.6 64.5 $\uparrow$3.1 74.8 $\uparrow$2.6 51.4 $\uparrow$0.3 60.3 $\uparrow$3.4 39.6 $\uparrow$2.6 40.3 $\uparrow$2.8 27.8 $\uparrow$3.0
I2T(Human) 85.7 $\uparrow$2.0 71.9 $\uparrow$4.4 82.0 $\uparrow$4.6 65.9 $\uparrow$4.5 77.6 $\uparrow$5.4 56.3 $\uparrow$5.2 63.3 $\uparrow$6.4 44.0 $\uparrow$6.9 45.6 $\uparrow$8.1 32.2 $\uparrow$7.4
Basic(RandAug) 84.1 $\uparrow$0.4 67.1 $\downarrow$0.4 78.7 $\uparrow$1.3 61.6 $\uparrow$0.2 70.9 $\downarrow$1.3 50.4 $\downarrow$0.7 57.8 $\uparrow$0.9 37.3 $\uparrow$0.2 35.6 $\downarrow$1.9 24.4 $\downarrow$0.4
T-I2I(SD) 83.8 $\uparrow$0.1 69.1 $\uparrow$1.6 77.7 $\uparrow$0.3 62.2 $\uparrow$0.8 75.3 $\uparrow$3.1 52.1 $\uparrow$1.0 59.7 $\uparrow$2.8 38.0 $\uparrow$1.0 39.3 $\uparrow$1.8 25.8 $\uparrow$1.0
T2I(SD) 83.3 $\downarrow$0.4 68.4 $\uparrow$0.9 76.1 $\downarrow$1.3 61.5 $\uparrow$0.1 73.3 $\uparrow$1.1 52.2 $\uparrow$1.1 57.2 $\uparrow$0.3 37.8 $\uparrow$0.7 37.9 $\uparrow$0.4 25.2 $\uparrow$0.4
Table 8: Comparison of CLIP trained on the Flickr30k dataset for ITR under the no-attack scenario (Clean), unimodal attacks (PGD and BERT), and multimodal attacks (Co-Attack and SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Method Img aug. Text. aug. Unimodal attacks Multimodal attacks
Clean PGD (image attack) BERT (text attack) Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1
Finetune 66.6 50.1 6.8 5.4 36.9 23.7 2.9 1.8 0.1 0.1
FARE 45.2 32.3 40.4 29.3 22.4 15.9 16.7 11.4 9.1 6.9
TeCoA-ITR 58.0 41.6 51.0 36.7 30.6 18.2 21.3 12.5 9.6 6.2
(ours) MAT 55.8 40.7 49.5 36.1 42.9 27.4 31.4 19.5 17.7 12.3
[…]
T2T(LangRW) 51.8 $\downarrow$4.0 37.7 $\downarrow$3.0 45.7 $\downarrow$3.8 33.5 $\downarrow$2.5 38.7 $\downarrow$4.2 26.1 $\downarrow$1.3 28.7 $\downarrow$2.7 19.0 $\downarrow$0.6 17.0 $\downarrow$0.7 12.6 $\uparrow$0.3
I2T(div-Caps) 56.7 $\uparrow$0.9 39.9 $\downarrow$0.8 51.2 $\uparrow$1.8 36.3 $\uparrow$0.2 40.2 $\downarrow$2.6 26.1 $\downarrow$1.3 30.3 $\downarrow$1.1 19.4 $\downarrow$0.2 18.9 $\uparrow$1.2 12.5 $\uparrow$0.2
I2T(Human) 58.9 $\uparrow$3.1 43.1 $\uparrow$2.4 52.2 $\uparrow$2.7 38.5 $\uparrow$2.4 44.7 $\uparrow$1.9 29.9 $\uparrow$2.4 33.6 $\uparrow$2.2 21.9 $\uparrow$2.3 21.3 $\uparrow$3.6 14.5 $\uparrow$2.2
Basic(RandAug) 55.9 $\uparrow$0.1 40.7 $\downarrow$0.0 49.4 $\downarrow$0.1 36.3 $\uparrow$0.2 41.7 $\downarrow$1.2 27.2 $\downarrow$0.2 30.7 $\downarrow$0.7 19.6 $\uparrow$0.1 18.3 $\uparrow$0.6 12.6 $\uparrow$0.3
T2I(SD) 54.2 $\downarrow$1.6 38.3 $\downarrow$2.4 46.3 $\downarrow$3.2 33.4 $\downarrow$2.7 39.5 $\downarrow$3.3 25.9 $\downarrow$1.5 28.5 $\downarrow$2.9 18.1 $\downarrow$1.5 15.9 $\downarrow$1.8 11.0 $\downarrow$1.3
T,I2I(SD) 57.1 $\uparrow$1.3 41.7 $\uparrow$1.0 49.7 $\uparrow$0.2 36.0 $\downarrow$0.1 41.9 $\downarrow$0.9 28.3 $\uparrow$0.8 31.3 $\downarrow$0.0 20.2 $\uparrow$0.7 18.8 $\uparrow$1.1 12.6 $\uparrow$0.2
     (N:N) T,I2I(SD) I2T(Human) 59.0 $\uparrow$3.2 43.3 $\uparrow$2.6 52.2 $\uparrow$2.7 38.0 $\uparrow$1.9 44.6 $\uparrow$1.7 29.8 $\uparrow$2.3 34.1 $\uparrow$2.8 21.2 $\uparrow$1.7 20.8 $\uparrow$3.1 14.2 $\uparrow$1.9
Table 9: Comparison of CLIP trained on the COCO dataset for ITR under the no-attack scenario (Clean), unimodal attacks (PGD and BERT), and multimodal attacks (Co-Attack and SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Method Img aug. Text. aug. Unimodal attacks Multimodal attacks
Clean PGD (image attack) BERT (text attack) Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1
Finetune 89.5 77.7 48.8 34.0 79.1 55.0 32.3 18.4 2.5 1.3
TeCoA-ITR 85.4 69.3 78.7 63.1 72.1 48.9 61.0 38.1 35.5 21.9
(ours) MAT 82.0 66.3 78.0 63.8 77.1 56.1 73.0 54.0 47.1 32.9
[…]
T2T(LangRW) 81.4 $\downarrow$0.6 67.5 $\uparrow$1.3 76.0 $\downarrow$2.0 62.5 $\downarrow$1.3 75.8 $\downarrow$1.3 56.2 $\uparrow$0.1 70.0 $\downarrow$3.0 51.1 $\downarrow$3.0 46.3 $\downarrow$0.8 32.4 $\downarrow$0.5
I2T(div-Caps) 85.6 $\uparrow$3.6 71.0 $\uparrow$4.8 81.7 $\uparrow$3.7 66.6 $\uparrow$2.8 79.7 $\uparrow$2.6 60.1 $\uparrow$3.9 75.6 $\uparrow$2.6 55.2 $\uparrow$1.2 48.8 $\uparrow$1.7 35.0 $\uparrow$2.1
I2T(Human) 85.8 $\uparrow$3.8 72.8 $\uparrow$6.5 80.2 $\uparrow$2.2 67.6 $\uparrow$3.8 81.5 $\uparrow$4.4 62.8 $\uparrow$6.7 77.5 $\uparrow$4.5 58.7 $\uparrow$4.7 52.9 $\uparrow$5.8 38.8 $\uparrow$5.9
Basic(RandAug) 82.2 $\uparrow$0.2 67.2 $\uparrow$0.9 78.3 $\uparrow$0.3 63.2 $\downarrow$0.6 76.9 $\downarrow$0.2 56.2 $\uparrow$0.1 72.7 $\downarrow$0.3 53.2 $\downarrow$0.8 48.3 $\uparrow$1.2 33.4 $\uparrow$0.5
T-I2I(SD) 85.1 $\uparrow$3.1 69.8 $\uparrow$3.6 81.7 $\uparrow$3.7 67.1 $\uparrow$3.3 79.4 $\uparrow$2.3 59.1 $\uparrow$2.9 74.9 $\uparrow$1.9 55.9 $\uparrow$1.8 55.2 $\uparrow$8.1 37.6 $\uparrow$4.8
T2I(SD) 83.7 $\uparrow$1.7 68.3 $\uparrow$2.1 79.5 $\uparrow$1.5 63.6 $\downarrow$0.2 76.7 $\downarrow$0.4 59.0 $\uparrow$2.9 72.4 $\downarrow$0.6 55.1 $\uparrow$1.1 52.0 $\uparrow$4.9 36.2 $\uparrow$3.3
Table 10: Comparison of ALBEF trained on the Flickr30k dataset for ITR under the no-attack scenario (Clean), unimodal attacks (PGD and BERT), and multimodal attacks (Co-Attack and SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Method Img aug. Text. aug. Unimodal attacks Multimodal attacks
Clean PGD (image attack) BERT (text attack) Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1
Finetune 69.9 53.6 30.0 17.4 49.0 31.4 17.2 9.1 1.0 0.7
TeCoA-ITR 64.8 48.6 49.5 38.1 48.5 28.5 38.2 19.4 14.2 9.5
(ours) MAT 63.9 46.2 60.2 38.3 55.4 36.9 52.9 32.1 31.2 21.2
[…]
T2T(LangRW) 59.7 $\downarrow$4.2 42.1 $\downarrow$4.1 46.6 $\downarrow$13.5 31.0 $\downarrow$7.3 51.4 $\downarrow$4.1 32.9 $\downarrow$4.0 34.6 $\downarrow$18.3 23.9 $\downarrow$8.2 25.4 $\downarrow$5.8 15.7 $\downarrow$5.5
I2T(div-Caps) 66.0 $\uparrow$2.0 49.9 $\uparrow$3.7 57.0 $\downarrow$3.2 37.9 $\downarrow$0.4 56.6 $\uparrow$1.2 38.5 $\uparrow$1.6 47.5 $\downarrow$5.4 26.4 $\downarrow$5.6 35.5 $\uparrow$4.3 20.3 $\downarrow$0.9
I2T(Human) 68.5 $\uparrow$4.5 49.1 $\uparrow$3.0 64.9 $\uparrow$4.7 37.3 $\downarrow$1.0 59.7 $\uparrow$4.3 39.9 $\uparrow$3.0 55.3 $\uparrow$2.4 29.3 $\downarrow$2.8 36.2 $\uparrow$4.9 23.5 $\uparrow$2.2
Basic(RandAug) 63.1 $\downarrow$0.9 48.3 $\uparrow$2.1 61.1 $\uparrow$0.9 46.0 $\uparrow$7.7 54.8 $\downarrow$0.7 38.3 $\uparrow$1.4 52.1 $\downarrow$0.8 35.7 $\uparrow$3.7 30.3 $\downarrow$1.0 21.7 $\uparrow$0.4
T2I(SD) 61.5 $\downarrow$2.4 46.0 $\downarrow$0.1 58.7 $\downarrow$1.5 37.8 $\downarrow$0.5 53.5 $\downarrow$2.0 37.5 $\uparrow$0.6 50.6 $\downarrow$2.3 33.0 $\uparrow$1.0 25.3 $\downarrow$5.9 18.6 $\downarrow$2.6
T,I2I(SD) 64.9 $\uparrow$0.9 47.1 $\uparrow$1.0 61.8 $\uparrow$1.6 36.1 $\downarrow$2.2 56.5 $\uparrow$1.1 38.1 $\uparrow$1.3 54.1 $\uparrow$1.2 29.2 $\downarrow$2.9 33.2 $\uparrow$2.0 22.4 $\uparrow$1.2
Table 11: Comparison of ALBEF trained on the COCO dataset for ITR under the no-attack scenario (Clean), unimodal attacks (PGD and BERT), and multimodal attacks (Co-Attack and SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Method Img aug. Text. aug. Unimodal attacks Multimodal attacks
Clean PGD (image attack) BERT (text attack) Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1
Finetune 72.9 57.5 50.9 35.7 39.2 27.5 18.4 9.9 1.2 1.1
TeCoA-ITR 64.6 51.8 63.8 50.3 37.6 24.9 31.2 20.2 20.2 13.9
(ours) MAT 66.9 49.9 67.8 49.5 54.7 37.1 50.5 32.8 31.3 21.0
[…]
T,I2I(SD) 68.2 $\uparrow$1.3 50.5 $\uparrow$0.6 62.3 $\downarrow$5.5 43.6 $\downarrow$6.0 55.0 $\uparrow$0.2 37.0 $\downarrow$0.0 47.9 $\downarrow$2.6 29.9 $\downarrow$3.0 33.5 $\uparrow$2.2 22.9 $\uparrow$1.9
Table 12: Comparison of BLIP trained on the COCO dataset for ITR under the no-attack scenario (Clean), unimodal attacks (PGD and BERT), and multimodal attacks (Co-Attack and SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
C.2 Visual grounding (VG)
Table 13 presents results of ALBEF trained on the COCO dataset for the VG task. The results highlight MAT’s effectiveness beyond ITR task, as well as the effectiveness of one-to-many augmentations. We observe that I2T(div-Caps) outperforms I2T(Human), which can be attributed to the design of the prompts used in I2T(div-Caps). Specifically, I2T(div-Caps) generates captions that focus on both foreground and background objects, providing a more comprehensive description of the image (see Sec. A.3 for the prompt design). This makes I2T(div-Caps) particularly useful for the VG task, where understanding diverse objects located in different areas of the image is crucial.
Method Img aug. Text. aug. Clean SGA
val test-A test-B val test-A test-B
Fine-tune 50.2 57.6 40.6 32.9 36.3 29.6
TeCoA-ITR 49.6 55.4 41.4 38.8 44.8 32.9
(ours) MAT 48.8 53.3 40.7 40.4 44.2 35.1
[…]
I2T(Human) 50.6 $\uparrow$1.8 55.9 $\uparrow$2.6 41.7 $\uparrow$1.1 42.0 $\uparrow$1.6 46.5 $\uparrow$2.3 35.7 $\uparrow$0.6
T,I2I(SD) 49.6 $\uparrow$0.7 55.9 $\uparrow$2.6 40.8 $\uparrow$0.1 41.1 $\uparrow$0.7 46.0 $\uparrow$1.8 34.8 $\downarrow$0.3
Table 13: Accuracy comparison of ALBEF trained on the COCO dataset for VG under the no-attack scenario (Clean) and multimodal attack (SGA).
Appendix D Qualitative Results
Finally, we present qualitative results for ITR. We compare the TeCoA-ITR baseline, a unimodal image defense method, with our proposed multimodal defense method, MAT+, which incorporates I2T(Human) augmentations. Both methods are evaluated against a multimodal attack (SGA). Figure 8 and Figure 9 illustrate the comparison for image retrieval and text retrieval results, respectively.
Figure 8: Qualitative comparison of image retrieval results under multimodal attack (SGA). We compare the TeCoA-ITR baseline (unimodal defense) with our proposed MAT+ (multimodal defense using I2T(Human) augmentations). Images with a blue border indicate correct retrieval, while those with a red border indicate incorrect retrieval.
Figure 9: Qualitative comparison of text retrieval results under multimodal attack (SGA). We compare the TeCoA-ITR baseline (unimodal defense) with our proposed MAT+ (multimodal defense using I2T(Human) augmentations).